# Bridging Domain Gap for Flight-Ready Spaceborne Vision

### v1

@@ 2.2.1 Source-Free Domain Adaptation
A different approach conducive to the operational constraints of space missions is source-free domain adaptation which does not require the availability of large-scale training images during the adaptation phase. Existing source-free algorithms leverage generative models for feature alignment [65, 66, 67] or pseudo-labeling and information maximization [68]. TENT [69] performs entropy minimization while updating only the affine parameters of the Batch Normalization (BN) layers [70]. However, methods based on entropy minimization require optimizations to be performed on batches of images in order to avoid trivial solutions, which could become computationally expensive on satellite avionics.
@@ 2.2.2 Spaceborne Applications
On the other hand, SPNv2 [18], one of the baseline models developed by the authors, designed a multi-scale, multi-task learning CNN architecture and trained with extensive data augmentations on the SPEED+ synthetic images only. The largest variant of SPNv2 would have ranked 3rd place in lightbox and 6th in sunlamp categories, respectively, without accessing the HIL domain images during training at all. Finally, EagerNet [80] performs dense predictions of pixel-wise object coordinates as opposed to heatmap predictions that many other methods adopted. They additionally predict the errors of predicted model coordinates which results in multiple pose hypotheses which are further refined using a probabilistic refinement. EagerNet trained with extensive data augmentation, including those specifically designed to target the HIL domain imagery, achieves state-of-the-art robustness without accessing the HIL domains, winning both HIL categories in the post-mortem competition.
[…]
@@ 3 Requirements of Spaceborne ML Models
Before introducing SPNv3, this section provides a brief overview of various requirements that must be met by not only SPNv3 but also ML models that are intended to run in space onboard satellite avionics. Some of these requirements, in addition to the OOD robustness across the sim2real gap, drive the design choices of SPNv3 in the next section.
@@ 3.1 Computationally Efficiency
As explained in section 1, it may be desirable to perform online training of NNs in space to fully close the domain gap by directly incorporating the flight images that only become available during RPOD [19]. In this case, the NN must also be computationally efficient to run not only forward but also backward gradient propagation. Unlike inference, the training most certainly requires a GPU for the operation. However, the bright side is that the training need not run for every acquired image but only when there has been a substantial “change” in the imagery compared to the previous training round. The motivation is to avoid overfitting the NN to the specific scenery (e.g., Earth background) and view of the target since they do not change so abruptly within the short time frame in space especially at higher altitudes. Therefore, as long as there is a GPU onboard, the training latency is likely to be less of a computational bottleneck than inference which should be running in real time.
3.1.1 Remark on Flight Heritage GPU
@@
Even nowadays, GPUs are still rarely found only in short-term missions which typically are not concerned with long-term radiation effects such as Total Ionising Dose (TID), or as part of a larger spacecraft which provides better protection of the onboard computing systems. Recent state-of-the-art report on small spacecraft technology by NASA has tabulated a number of GPU-based avionics [83, Table 8.1, §8.3], but many Commercially Off-The-Shelf (COTS) systems such as NVIDIA Jetson series are still not flight-proven or are radiation tolerant only against Single Event Effects (SEE) (e.g., bit flips).
@@
However, there is an increasing interest in deploying GPUs into deep space or long-term missions in LEO with existing flight heritage [25]. For example, Aitech Systems, Inc., a company that manufactures rugged computers for military and aerospace applications, utilizes the NVIDIA Jetson TX2i System-on-Module (SoM) in their S-A1760 system [24] which is designed for spacecraft and small satellites and has a flight heritage in LEO as part of a larger payload system [84]. The startup company Aethero [^6] https://www.aethero.com/ is also developing radiation-hardened edge computers for on-orbit data processing and autonomous decision-making in space. Its ECM-NxN utilizes an NVIDIA Jetson Orin processor, considered the best GPU edge processor available. While the ECM-NxN module has not directly flown in space, Aethero’s use of this processor in their space computer indicates potential applications in space exploration. Finally, the Ingenuity Mars Helicopter used a GPU-equipped Qualcomm® Snapdragon 801 processor with great results, even though it operated in a “terrestrial” environment [85]. These are a few examples of global efforts to deploy edge GPU systems to space, and there is no doubt that more and more powerful space-grade GPUs will become commercially available to endure the harsh space environment for a prolonged period of time.
@@ 3.2 Number of Parameters
The number of parameters is closely related to the computational efficiency since more weights translate to more operations to be performed for both training and inference. However, the NN size is also tightly connected to the power consumption. For instance, the seminal work of Han et al. [86] in 2015 identified that, under the 45nm Complementary Metal-Oxide Semiconductor (CMOS) technology, 32-bit memory access to a Dynamic Random Access Memory (DRAM) consumes three orders of magnitude more power than accessing a Static Random Access Memory (SRAM) cache or performing a 32-bit floating point arithmetic. Considering miniature satellite systems such as CubeSats with limited power generation capability, it would be favorable to minimize the NN size so that the entire network can be hosted on a local SRAM which is limited in memory size (typically up to 20MB for radiation hardened SRAM [83, Table 8.2, §8.3.4]).
@@ 3.3 Batch-Agnostic Architecture
where $\bm{X}_{i}$ is the feature map at the $i$-th layer. Unlike other methods, BN aims to compute the batch-wise feature map statistics across the entire training set. Since this is infeasible, in practice, the NORM layer statistics are computed across mini-batches, and the “global” statistics are updated using a running average as training progresses. Then, during inference, these running averages of the mean and variance are fixed and used for normalization.
@@
The problem is that these approximations of the training data distribution in BN layers are simply incompatible with those of the test data that are drawn from a different distribution. Therefore, in this work, two methods are considered in search of a robust and batch-agnostic pose estimation NN architecture. One is to simply replace the BN layers in existing pre-trained architectures with batch-agnostic NORM layers such as Layer Normalization (LN) [93] and Group Normalization (GN) [94]. The other is to design or adopt a NN architecture that are inherently without BN layers such as ViT [22].
4 Searching for Next Spacecraft Pose Network (SPNv3)
The goal of this section is two-fold. One is to introduce the overall pose estimation architecture given an image of the target at a visible distance in an RPOD scenario (section 4.1). Then, different NN architectures (section 4.2) and data augmentation techniques (section 4.3) for training a pose estimation NN are explained which are combined in various configurations to study and explore which aspects of NN architecture and training algorithm contribute the most to the OOD robustness and reduced latency of a pose estimation NN (Section 5). While SPN is the name of the first spacecraft pose estimation CNN introduced by Sharma and D’Amico [6], it is also used throughout this work as a legacy name referring to all NN architectures designed or adopted for spacecraft pose estimation by SLAB. When referring to the first original SPN, it is always accompanied by the reference to the relevant work [6].
4.1 Operational Scenario
Figure 3: Visualization of far-range AON [95] (left); close-range pose estimation on a cropped image (middle); and mechanism of feature/heatmap extraction by a NN (right). In general, the sensors are selected to ensure overlapping working distance to the target but it is not always possible based on the orbit and size of the target.
@@ 4.1 Operational Scenario
In all vision-based RPOD scenarios, the servicer spacecraft begins tracking the non-cooperative target at kilometers of separations as shown in Fig. 3. For example, Angles-Only Navigation (AON) [96, 97, 95, 98] obtains bearing angle measurements of the target from a Narrow Field-Of-View (NFOV) camera such as a star tracker which allows tracking of the target’s relative orbital state via nonlinear filtering. AON continues until the inter-spacecraft distance becomes small enough such that the target appears resolved in the camera view, at which point the pose estimation algorithm kicks in.
@@
Given the scenario, this work adopts the two-stage approach visualized in Fig. 3 as the common mechanism of the ML-based pose estimation algorithm which was first independently proposed by Park et al. [32] and Chen et al. [33] for SPEC2019 [31]:
@@
[…]
@@ 4.2 Pose Estimation Architectures
Figure 4: Visualization of different pose estimation architectures: EfficientDet [100], HRNet [101] and ViTPose [102].
This section explores three different heatmap detection architectures—EfficientDet [100], HRNet [101] and ViTPose [102]—visualized in Fig. 4. Note that this is not a comprehensive list of state-of-the-art pose estimation NN architectures. Nonetheless, they are chosen due to their simplicity and in order to keep the number of training sessions tractable while exploring vastly different architectural options.
@@ 4.2.1 EfficientDet
[…]
@@ 4.2.2 HRNet
High Resolution Network (HRNet) [101] is also designed for performing tasks that require multi-scale predictions. Recall that EfficientDet builds upon a CNN backbone that is typically designed for the image classification task, resulting in an architecture which continuously reduces the feature resolution all the way down to 1/32 of the original input. On the other hand, HRNet is designed to maintain high-resolution features throughout the entire forward propagation, outputting the multi-resolution features by the nature of its design. To facilitate learning at different scales, HRNet first processes the input image to 1/4 resolution, then gradually adds lower-resolution features in parallel, continuously fusing information across different scales as visualized in Fig. 4.
@@ 4.3 Extensive Data Augmentation
Data augmentation is vital to training NNs that generalize well beyond their training data distribution and to preventing overfitting. As basic sets of augmentation, SPN heavily leverages the Albumentations library [104] for various image processing operations. For each training sample, $N$ augmentation operations are randomly chosen from the set of designated operations in a manner similar to RandAugment [105]. Specifically, while RandAugment utilizes a single hyperparameter to control the constant level of magnitude of these operations (e.g., standard deviation of white Gaussian noise), this work simply picks each operation with a random magnitude within some pre-defined window. Then, these $N$ operations are applied to the original image in sequence. The complete list of employed augmentations is provided in Table 1 along with their equivalent Albumentations commands. All commands are run with the default range of magnitudes.
@@
In addition to the RandAugment-like augmentations which are composed of simple image processing operations such as Gaussian noise and contrast adjustment, three additional operations are explored that were designed specifically to address the domain gap issue.
@@ 4.3.1 Style Augmentation
Geirhos et al. [63] has empirically shown in 2019 that many state-of-the-art CNN architectures, such as ResNet [107] and VGG [108], exhibit strong bias toward object’s texture when trained on ImageNet [34]. This finding is very counter-intuitive for humans who would, for example, classify an object or an animal based on its global or a set of local shapes, not based on its texture. To demonstrate such behavior, the authors have created Stylized-ImageNet (SIN) dataset by applying the AdaIN Neural Style Transfer (NST) pipeline [109] to each image in ImageNet with random styles from Kaggle’s Painter by Numbers dataset [^7] https://www.kaggle.com/c/painter-by-numbers/. The result shows that when the same CNN architectures are trained on the SIN dataset, the networks not only exceed the performance of those trained on ImageNet, but they also show human-level robustness to previously unseen image distortions, such as noise, contrast change, and high- or low-pass filtering.
@@ 5 Trade-Off Analyses: OOD Robustness and Latency
This section performs the trade-off analyses between robustness across domain gap and training/inference latency for various configurations of pose estimation architectures (section 4.2), data augmentation techniques (section 4.3), and other aspects of NN training such as transfer learning and input resolution. The ultimate goal is to find the configuration that is both robust on SPEED+ HIL domain images and computationally efficient on a representative hardware, capable of operating above nominal GN&C update frequency (section 3.1) while also being batch-agnostic for potential online training (section 3.3).

### v2

@@ 2.2.1 Source-Free Domain Adaptation
A different approach conducive to the operational constraints of space missions is source-free domain adaptation, which does not require the availability of large-scale training images during the adaptation phase. Existing source-free algorithms leverage generative models for feature alignment [71, 72, 73] or pseudo-labeling and information maximization [74]. TENT [75] performs entropy minimization while updating only the affine parameters of the Batch Normalization (BN) layers [76]. However, methods based on entropy minimization require optimizations to be performed on batches of images in order to avoid trivial solutions, which could become computationally expensive on satellite avionics.
@@ 2.2.2 Spaceborne Applications
On the other hand, SPNv2 [22], one of the baseline models developed by the authors, designed a multi-scale, multi-task learning CNN architecture and trained with extensive data augmentations on the SPEED+ synthetic images only. The largest variant of SPNv2 would have ranked 3rd place in lightbox and 6th in sunlamp categories, respectively, without accessing the HIL domain images during training at all. Finally, EagerNet [86] performs dense predictions of pixel-wise object coordinates as opposed to heatmap predictions that many other methods adopted. They additionally predict the errors of predicted model coordinates, which results in multiple pose hypotheses that are further refined using a probabilistic refinement. EagerNet trained with extensive data augmentation, including those specifically designed to target the HIL domain imagery, achieves state-of-the-art robustness without accessing the HIL domains, winning both HIL categories in the post-mortem competition.
As shown later, SPNv3 is able to achieve comparable robustness relative to EagerNet without access to the HIL domains as well. However, SPNv3 is a much simpler architecture that outputs heatmaps about known keypoints on the target surfaces, which can be provided directly to the onboard navigation filter [20] or be used to solve for the optimal pose solution via PnnP [87].
@@ 3 Requirements of Spaceborne ML Models
Before introducing SPNv3, this section provides a brief overview of various requirements that must be met by not only SPNv3, but also ML models that are intended to run in space onboard satellite avionics. Some of these requirements, in addition to the OOD robustness across the sim2real gap, drive the design choices of SPNv3 in the next section.
@@ 3.1 Computationally Efficiency
As explained in section˜1, it may be desirable to perform additional online training of NNs in space to fully close the domain gap by directly incorporating the flight images that only become available during RPOD. In this case, the NN must also be computationally efficient to run not only forward but also backward gradient propagation. Unlike inference, the training most certainly requires a GPU as it is computationally expensive to track gradients across the entire NN. However, Park and D’Amico [23] has shown that performing one training iteration whenever there has been a substantial change in the spacecraft pose is sufficient to improve the NN’s OOD robustness. This not only renders onboard training computationally feasible but also prevents overfitting the NN to the specific scenery (e.g., Earth background) and view of the target. Therefore, under this strategy, as long as there is a GPU onboard, the training latency is likely to be less of a computational bottleneck than inference, which should be running in real-time. Readers are referred to Park and D’Amico [23] for more details on online supervised training of NN onboard spacecraft avionics.
3.1.1 Remark on GPU Flight Heritage
@@
Even nowadays, GPUs are still rarely found onboard spacecraft, and only in short-term missions which typically are not concerned with long-term radiation effects such as Total Ionising Dose (TID), or as part of a larger spacecraft which provides better protection of the onboard computing systems. Recent state-of-the-art report on small spacecraft technology by NASA has tabulated a number of GPU-based avionics [89, Table 8.1, §8.3], but many Commercially Off-The-Shelf (COTS) systems such as NVIDIA Jetson series are still not flight-proven or are radiation tolerant only against Single Event Effects (SEE) (e.g., bit flips).
@@
However, there is an increasing interest in deploying GPUs into deep space or long-term missions in LEO with existing flight heritage [29]. For example, Aitech Systems, Inc., a company that manufactures rugged computers for military and aerospace applications, utilizes the NVIDIA Jetson TX2i System-on-Module (SoM) in their S-A1760 system [28] which is designed for spacecraft and small satellites and has a flight heritage in LEO as part of a larger payload system [90]. The startup company Aethero [^6] https://www.aethero.com/ is also developing radiation-hardened edge computers for on-orbit data processing and autonomous decision-making in space. Its ECM-NxN utilizes an NVIDIA Jetson Orin processor, considered the best GPU edge processor available. While the ECM-NxN module has not directly flown in space, Aethero’s use of this processor in their space computer indicates potential applications in space exploration. Finally, the Ingenuity Mars Helicopter used a GPU-equipped Qualcomm® Snapdragon 801 processor with great results, even though it operated in a “planetary” environment [91]. These are a few examples of global efforts to deploy edge GPU systems to space, and there is no doubt that more and more powerful space-grade GPUs will become commercially available to endure the harsh space environment for a prolonged period of time.
@@ 3.2 Number of Parameters
The number of NN parameters is closely related to the computational efficiency since more weights translate to more operations to be performed for both training and inference. However, the NN size is also tightly connected to the power consumption. For instance, the seminal work of Han et al. [92] in 2015 identified that, under the 45nm Complementary Metal-Oxide Semiconductor (CMOS) technology, 32-bit memory access to a Dynamic Random Access Memory (DRAM) consumes three orders of magnitude more power than accessing a Static Random Access Memory (SRAM) cache or performing a 32-bit floating point arithmetic. Considering miniature satellite systems such as CubeSats with limited power generation capability, it would be favorable to minimize the NN size so that the entire network can be hosted on a local SRAM, which is limited in memory size (typically up to 20MB for radiation hardened SRAM [89, Table 8.2, §8.3.4]).
@@ 3.3 Batch-Agnostic Architecture
where $\bm{X}_{i}$ is the feature map at the $i$-th layer. Unlike other methods, BN aims to compute the batch-wise feature map statistics across the entire training set. However, computing such statistics across the entire training domain is practically infeasible given a limited dataset. Therefore, the BN layer statistics are instead computed across mini-batches, and the approximate “global” statistics are updated using a running average as training progresses. Then, during inference, these running averages of the mean and variance are fixed and used for normalization.
@@
The problem is that these approximations of the training data distribution in BN layers are simply incompatible with those of the test data that are drawn from a different distribution. Therefore, in this work, two methods are considered in search of a robust and batch-agnostic pose estimation NN architecture. One is to simply replace the BN layers in existing pre-trained architectures with batch-agnostic NORM layers such as Layer Normalization (LN) [99] and Group Normalization (GN) [100]. The other is to design or adopt a NN architecture that is inherently without BN layers such as ViT [26].
3.4 Summary
In summary, in order to ensure that SPNv3 is suitable for real-time onboard inference and fine-tuning, it must meet the following objectives:
  • it must be able to process incoming images within a representative ADCS system frequency (e.g., 0.5 Hz)
  • it must have a minimum number of learnable parameters in order to reduce the memory footage
  • considering the possibility of onboard training and fine-tuning, it must have minimum dependence on NN training components that must operate on a batch of images (e.g., BN layers)
All the above must be accomplished while simultaneously maximizing the pose estimation accuracy on OOD spaceborne images.
4 Pose Estimation Pipeline
Figure 3: Visualization of far-range AON [101] (left) and close-range pose estimation (right).
@@ 4 Pose Estimation Pipeline
In all vision-based RPOD scenarios, the servicer spacecraft begins tracking the non-cooperative target at kilometers of separations as shown in Fig.˜3. For example, Angles-Only Navigation (AON) [102, 103, 101, 104] obtains bearing angle measurements of the target from a Narrow Field-Of-View (NFOV) camera, such as a star tracker, which allows tracking of the target’s relative orbital state via nonlinear filtering. AON continues until the inter-spacecraft distance becomes small enough such that the target appears resolved in the camera view, at which point the pose estimation algorithm kicks in. Note that in general, the sensors are selected to ensure overlapping working distance to the target during the transition from far-range to close-range RPO, but it is not always possible based on the orbit and size of the target.
Figure 4: Visualization of the pose estimation pipeline adopted in this work.
@@
Given the scenario, this work adopts the three-stage approach visualized in Fig.˜4 as the common mechanism of the ML-based pose estimation algorithm, which was first independently proposed by Park et al. [37] and Chen et al. [38] for SPEC2019 [36]:
@@
    The second stage is the main pose estimation NN, which takes the cropped low-resolution images and outputs $K$ heatmaps centered at the 2D locations of known target surface keypoints. The advantage of dense predictions such as heatmaps is that their spread about the peak can be interpreted as uncertainties associated with each prediction [43]. Therefore, at each $k$-th heatmap, in addition to its mean location $(x_{k},y_{k})$, its spread is approximated as a covariance matrix $\bm{\Sigma}_{k}$ and returned.
  3. 3.
    The final stage takes the 2D keypoint locations and, knowing the 3D coordinates of these keypoints in the body frame, solves PnnP to compute the full 6D pose solution. The uncertainty covariances can also be incorporated into the uncertainty-aware PnnP algorithm [105, 43], or the mean and covariance can both be provided as measurements to the navigation filter [20] which tracks the target’s pose.
@@
5 Searching for Next Spacecraft Pose Network (SPNv3)
In this section, different NN architectures (section˜5.1) and data augmentation techniques (section˜5.2) for training a pose estimation NN are explained which are combined in various configurations to study and explore which aspects of NN architecture and training algorithm contribute the most to the OOD robustness and reduced latency of a pose estimation NN (Section 6). While SPN is the name of the first spacecraft pose estimation CNN introduced by Sharma and D’Amico [10], it is also used throughout this work as a legacy name referring to all NN architectures designed or adopted for spacecraft pose estimation by SLAB. When referring to the first original SPN, it is always accompanied by the reference to the relevant work [10].
@@ 5.1 Pose Estimation Architectures
Figure 5: Visualization of different pose estimation architectures: EfficientDet [106], HRNet [107] and ViTPose [108]. The architectures commonly consist of a backbone/encoder network and a heatmap prediction head.
This section explores three different heatmap detection architectures—EfficientDet [106], HRNet [107] and ViTPose [108]—visualized in Fig.˜5. Note that this is not a comprehensive list of state-of-the-art pose estimation NN architectures. Nonetheless, they are chosen due to their simplicity and in order to keep the number of training sessions tractable while exploring vastly different architectural options.
@@ 5.1.1 EfficientDet
EfficientDet was originally proposed by Tan et al. [106] to enable computationally efficient yet high-performance object detection and semantic segmentation. EfficientDet builds upon the EfficientNet [24] backbone designed for the image classification task. At the heart of EfficientDet is the Bidirectional Feature Pyramid Network (BiFPN), a novel FPN architecture [109] that fuses and processes feature outputs from the backbone at multiple scales. By fusing features from low- to high-resolution and vice versa, BiFPN outperforms previous state-of-the-art FPN architectures on various benchmark object detection and semantic segmentation datasets. In the end, the feature outputs from BiFPN are processed by the task-specific prediction heads composed of convolution layers. For tasks that require multi-scale predictions (e.g., object detection), features from all resolutions are processed. In this work, the heatmap prediction head is attached only to the P2 level feature as shown in Fig.˜5, where Pnn indicates that the feature size is 1 / $2^{n}$ of the input image resolution. Only the P2 features (i.e., heatmap has 1/4 of the input resolution) are used since the accuracy of heatmap prediction heavily depends on the heatmap pixel resolution.
@@ 5.1.2 HRNet
High Resolution Network (HRNet) [107] is also designed for performing tasks that require multi-scale predictions. Recall that EfficientDet builds upon a CNN backbone that is typically designed for the image classification task, resulting in an architecture which continuously reduces the feature resolution all the way down to 1/32 of the original input. On the other hand, HRNet is designed to maintain high-resolution features throughout the entire forward propagation, outputting the multi-resolution features by the nature of its design. To facilitate learning at different scales, HRNet first processes the input image to 1/4 resolution, then gradually adds lower-resolution features in parallel, continuously fusing information across different scales as visualized in Fig.˜5.
@@ 5.2 Extensive Data Augmentation
Data augmentation is vital to training NNs that generalize well beyond their training data distribution and to preventing overfitting. As basic sets of augmentation, SPN heavily leverages the Albumentations library [110] for various image processing operations. For each training sample, $N$ augmentation operations are randomly chosen from the set of designated operations in a manner similar to RandAugment [111]. Specifically, while RandAugment utilizes a single hyperparameter to control the constant level of magnitude of these operations (e.g., standard deviation of white Gaussian noise), this work simply picks each operation with a random magnitude within some pre-defined window. Then, these $N$ operations are applied to the original image in sequence. The complete list of employed augmentations is provided in Table˜1 along with their equivalent Albumentations commands. All commands are run with the default range of magnitudes.
@@
In addition to the RandAugment-like augmentations which are composed of simple image processing operations such as Gaussian noise and contrast adjustment, three additional operations are explored that were designed specifically to address the domain gap issue. The effects of these operations are visualized in Fig.˜6.
@@ 5.2.1 Style Augmentation
Geirhos et al. [69] has empirically shown in 2019 that many state-of-the-art CNN architectures, such as ResNet [113] and VGG [114], exhibit strong bias toward object’s texture when trained on ImageNet [39]. This finding is very counter-intuitive for humans who would, for example, classify an object or an animal based on its global or a set of local shapes, not based on its texture. To demonstrate such behavior, the authors have created the Stylized-ImageNet (SIN) dataset by applying the AdaIN Neural Style Transfer (NST) pipeline [115] to each image in ImageNet with random styles from Kaggle’s Painter by Numbers dataset [^7] https://www.kaggle.com/c/painter-by-numbers/. The result shows that when the same CNN architectures are trained on the SIN dataset, the networks not only exceed the performance of those trained on ImageNet, but they also show human-level robustness to previously unseen image distortions, such as noise, contrast change, and high- or low-pass filtering.
@@ 6 Trade-Off Analyses: OOD Robustness and Latency
This section performs the trade-off analyses between robustness across domain gap and training/inference latency for various configurations of pose estimation architectures (section˜5.1), data augmentation techniques (section˜5.2), and other aspects of NN training such as transfer learning and input resolution. The ultimate goal is to find the configuration that is both robust on SPEED+ HIL domain images and computationally efficient on a representative hardware, capable of operating above nominal GN&C update frequency (section˜3.1) while also being batch-agnostic for potential online training (section˜3.3).